OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates ordered pair negative 7 comma 5 pi divided by 3. ordered pair 7 divided by 2 comma negative 7 divided by 2 ordered pair 7 square root 3 divided by 2 comma negative 7 divided by 2 ordered pair negative 7 divided by 2 comma 7 divided by 2 ordered pair negative 7 divided by 2 comma 7 square root 3 divided by 2

OpenStudy (anonymous):

@beccaboo333

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I think the answer is the last one

ganeshie8 (ganeshie8):

\(\large x = r\cos \theta\) \(\large y = r\sin\theta\)

OpenStudy (anonymous):

So would it be the last one?

OpenStudy (anonymous):

This is what it looks like

ganeshie8 (ganeshie8):

\(\large x = r\cos \theta = -7*\cos(5 \pi / 3) = \frac{-7}{2}\)

OpenStudy (anonymous):

and then after that comes the 7sqrt3/2 right? :)

ganeshie8 (ganeshie8):

\(\large y = r\sin \theta = -7*\sin(5 \pi / 3) = \frac{7\sqrt{3}}{2}\)

OpenStudy (anonymous):

yay! t hanku

ganeshie8 (ganeshie8):

yes :) good job !